# MAROON: A Framework for the Joint Characterization of Near-Field High-Resolution Radar and Optical Depth Imaging Techniques

### v2

@@ 9. Discussion of MIMO Radar Signal and Reconstruction
Having found no direct relation between signal magnitude and the reconstruction quality, we now examine the depth deviation concerning the previously mentioned three object quantities, which are shown in the bottom row of Figure 6. Similar to our findings for signal magnitude, metal objects generally exhibit the lowest depth deviation with a relatively small spread compared to other material classes, indicating that object material influences reconstruction quality. Furthermore, we find no direct relationship between object size and reconstruction quality. The most notable trend is seen with varying object geometry, where depth error increases alongside the median surface incidence angle across all material classes. The backprojection algorithm assumes that similar energy amounts are received at a point across the majority of RX antennas. Received energy diminishes for surfaces oriented away from the antenna geometry, leading to variations based on antenna positioning. This, in turn, can lead to reduced confidence in the measurements, causing valid data to be filtered out along with noise.
To conclude, the analysis of MIMO imaging radar reconstructions based on the received signal response is considerably less straightforward than that of optical sensors. Our findings suggest that the reconstruction quality of RF ToF sensors is primarily influenced by object geometry, followed by the influence of object material, especially in depth filtering with respect to signal confidence.
9.1. Limitations
Even though the depth deviation of the MIMO imaging radar is on par with that of optical sensors, the reconstructions exhibit significantly more holes, where no valid depth is estimated. While it is intuitive to assume that reconstruction quality is influenced by object material (limiting the returned signal amount — akin to optically transmissive materials), we observe that in our experiments the object geometry is the primary factor of influence and has a greater impact than in optical sensors; however, disentangling the effects of geometry and material remains challenging, as precise impacts on fine-grained surface details cannot be easily assessed. Without a direct mapping between point targets and RX antennas, depth evaluation concerning these surface-level details is infeasible without backprojection. On the other hand, the reconstructed outcomes of backprojection may not align with reality, as the method incorporates a systematic bias by relying on the Born approximation (Ahmed, 2014). To overcome these limitations, we suggest using simulation frameworks, such as employed by Schuessler et al. (2021) or Bräunig et al. (2023a). These could help disentangle potential error sources behind missing reconstructions and enhance the current signal processing methods.
Furthermore, it should be noted that the object reconstructions were evaluated solely for valid locations in the ground-truth data, excluding artifacts like ghost targets or other forms of noise that may arise from violations of the Born approximation. Lastly, we did not capture different orientations of flat objects, which would be an interesting future direction to investigate object orientation in isolation from geometry complexity.
10. Conclusion
In this paper, we present a novel multimodal dataset, MAROON, that allows us to characterize, for the first time, near-field MIMO imaging radars in direct relation with traditional depth imagers from the optical frequency domain for close-range applications. The dataset comprises depth images of a variety of objects, synchronously captured by four mutually calibrated depth imagers and a ground-truth multi-view stereo system. We subsequently analyzed the data within a comprehensive evaluation framework, offering quantitative and qualitative perspectives on each sensor’s depth deviation across multiple metric types, objects, and object-to-sensor distances.
In summary, our analysis revealed the following insights:
@@ 10. Conclusion
We hope that the public release of the MAROON dataset will give rise to further study of multimodal sensors in a joint reference frame. The findings presented here are based on aggregate trends and individual object analyses that contribute to the understanding of the addressed sensor characteristics; however, we believe that our dataset still invites further analysis, exploiting the high diversity of the 45 objects that could not be fully addressed in the scope of this paper. Furthermore, we anticipate that insights into the complementary strengths of optical and near-field radio-frequency depth imagers will benefit future sensor fusion applications.

### v4

@@ 7.3.2. Radio-Frequency Depth Deviation
While we find no direct relationship of object size to depth deviation, the most notable trend is seen with varying object geometry, where the deviation increases alongside the median surface incidence angle across all material classes. The backprojection algorithm assumes that similar energy amounts are received at a point across the majority of RX antennas. Received energy diminishes for surfaces oriented away from the antenna geometry, leading to variations based on antenna positioning. This, in turn, can lead to reduced confidence in the measurements, causing valid data to be filtered out along with noise.
Note that the MIMO radar we use has a large aperture and a high number of RX-TX antenna pairs, suggesting that stronger orientation-dependent effects may be observed with typical lower-resolution devices. To explore this further, we simulated various down-sampled antenna architectures and present their respective depth deviation in comparison to the fully occupied antenna array in the supplementary material.
7.3.3. Relation between RF Signal Response and Depth Deviation
Figure 5 visualizes potential correlations between the RF signal response and depth deviation in the right scatter plot. Focusing on the most prevalent material groups of polymer and metal objects, we generally find no direct relationship between signal magnitude and depth deviation. Polymer and metal objects have a large spread in the $x$- and $y$-axis, respectively, while the opposite axis has comparably low variation. The received signal magnitude may have a more significant influence on the reconstruction quality in less constrained scenarios, involving multiple objects and signal sources, where, for example, depth filtering becomes increasingly relevant. In our experiments, within the RF near field, we suggest that reconstruction quality is more closely related to the distribution of received signals across antennas, influenced primarily by the local antenna layout and characteristics.
To conclude, the analysis of MIMO imaging radar reconstruction quality is not as straightforward as that of optical sensors. We find no direct relationship between the RF signal response and the corresponding depth deviation after reconstruction and filtering on a global scale. Consistent patterns within object subsets suggest that the reconstruction quality of RF ToF sensors is primarily influenced by object geometry, while the impact of object material should not be overlooked, as it is a crucial factor for depth filtering.
To disentangle material and geometry-dependent effects at surface level, we suggest that material characterization is the first essential step in addressing this challenge; we will pick up on this topic in the subsequent application section.
8. Applications of MAROON
In this section, we highlight two applications utilizing the data from our proposed dataset. First, we briefly summarize the insights revealed from previous experiments:
@@ 8. Applications of MAROON
Given these insights, we anticipate that multimodal depth sensing amplifies the sensors’ complementary strengths, hence providing notable benefits for close-range applications — similar to the multi-sensor design employed in far-range sensing for self-driving cars.
We provide two examples, where we build upon prior work and demonstrate the substantial role of the high-quality sensor co-localizations from MAROON for their successful implementation: first, we show how the dataset is utilized for achieving realistic RF simulations by adapting and extending the concurrent research of Hofmann et al. (2025). Leveraging a pre-release of our dataset, Hofmann et al. propose a differentiable ray tracing pipeline to determine the material parameters of our captured objects.
Second, we present extended experiments for a recently proposed multimodal high-speed radar reconstruction method, MM-2FSK (Wirth et al., 2025), which utilizes only two frequencies as opposed to our employed 128 frequency-stepped backprojection. We note that we also examine the depth deviation in less computationally intensive versions of backprojection, by varying the frequency configuration, as detailed in the ablation study included in the supplementary material.
MVS Data Result 100% Target 100% Result 50% Target 50% Result 25% Target 25%
Cardb.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Plunger
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Hand Pr.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S1 Hand
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Disk
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Plate
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7. Inverse radar rendering results for different antenna configuration. 100% considers all transmitting and receiving antennas in the MIMO array ($94\times 94$), 50% considers every second respective antenna in the array ($47\times 47$), and 25% only considers every fourth antenna in the array ($24\times 24$). Despite obvious artifacts and ambiguities due to a reduced antenna density, the optimization process still performs robustly, and thus generalizes to other MIMO configurations. Objects from top to bottom: Cardboard, Plunger, Hand Printed F, S1 Hand Open Metal Disk (Thin), and V1 Metal Plate.
8.1. Material Characterization with Differentiable Ray Tracing
Accurate material modeling is vital not only for isolating the effects of mmWave signal interaction but also for high-fidelity radar simulation. Highly reflective materials, such as metals, produce vastly different radar returns compared to largely diffuse scatterers, such as objects made of wood or rubber. Therefore, Hofmann et al. (2025) propose a data-driven approach to determine reflective properties under mmWave radiation, such as permittivity and permeability.
They initialize their differentiable optimization pipeline by simulating radar returns with randomized material properties and compare the result to the real RF ToF sensor data from MAROON, while utilizing the MVS data as ground-truth geometry. Analogous to neural networks, where parameters are iteratively optimized using gradient descent, they continuously update the material properties of the object until the difference between the simulated and measured radar returns is minimal. Notably, the loss is computed on the raw phasor data, instead of reconstructed images, which increased both robustness and fidelity of the optimization due to bypassing artifacts introduced by the reconstruction algorithm. To facilitate a close match between the simulated and real phasor data, the radar gain and a small registration offset of $\frac{\lambda}{2}$ along each principal axis, where $\lambda$ is the longest wavelength emitted in a FSCW sequence, is optimized alongside the material properties (Hofmann et al., 2025). This registration offset requires the error in the MVS data to be smaller than one wavelength to avoid getting stuck in local minima due to ambiguities from recurring wave patterns. Fortunately, we can safely assume this to be the case in MAROON with the calibration error ranging from 1–2 mm, as discussed in Section 5.2, which is half of the mean wavelength of $\approx$ 4 mm in the worst case.
To demonstrate the versatility of the dataset, in addition to the original experiments conducted by Hofmann et al. (2025), we utilized the high number of antenna signals available in MAROON to examine the impact of a varying antenna configurations and aperture sizes. To this end, we used 100%, 50%, and 25% of the antennas, which were simulated by selecting every, every second, or every fourth RX/TX antenna from the raw phasor data in the dataset, respectively. We showcase results for the three different antenna configurations and six different objects in Figure 7.
For a visualization of the respective antenna apertures, we refer to the supplementary material, where we also conducted ablation studies of more consumer-friendly aperture configurations.
8.2. Multimodal Depth Sensing
While the backprojection algorithm is employed in many near-field high-resolution RF ToF applications, its reconstruction time is typically orders of magnitude slower than the sensor’s capture rate (see Table 1). As a robust and computationally efficient alternative, Wirth et al. (2025) introduce a multimodal image reconstruction method that builds upon the previous frequency shift keying (FSK) approach proposed by Bräunig et al. (2023). By utilizing an optical depth camera as a secondary sensor, point-wise depth priors are integrated into the 2FSK signal processing pipeline, allowing just two frequencies to adjust these depth priors towards the target object’s actual depth. The depth prior is essential to determine the correct period of the sinusoidal wave signal that is otherwise limited to a small unambiguous range. We refer the interested reader to (Wirth et al., 2025) for all technical details about the algorithm. The authors evaluated the proposed MM-2FSK method using the active stereo depth sensor and MIMO imaging radar from our pre-released dataset. In this section, we extend their work by comparing the method across all optical sensors in our dataset, simulating various capture scenarios influenced by the optical depth sensor.
Ablation with respect to Optical Depth Sensors.
Drawing from insights about sensor-specific characteristics, we examine how different depth imagers affect depth deviations in the MM-2FSK method. We follow the evaluation procedure detailed in (Wirth et al., 2025), employing the most promising frequency configuration — specifically, two frequencies at 72 and 82 GHz, resulting in a frequency difference of $\Delta f=10$ GHz.
In Figure 8, we display top-down views of the MM-2FSK reconstructions for three objects, overlaid with the ground-truth point cloud while varying the sensor that provides the depth prior. For the Flowerpot (Transparent) (top row), we notice numerous points reconstructed behind the object for all depth imagers except the ground truth. This transparency causes parts of the background or ground surface to be reconstructed, resulting in a depth prior positioned behind the ground truth. With limited unambiguous depth correction capabilities (Wirth et al., 2025), the MM-2FSK method can not correct outliers when the optical depth prior lies within a different signal period than the ground truth.
Similarly, for the V2 Metal Plate, large areas of its surface are reconstructed behind the object for both NIR ToF and active stereo sensors, due to multi-path effects and sensor oversaturation arising from the object’s perfect specularity.
Lastly, there are a few depth outliers behind the Bunny, primarily associated with the NIR ToF and active stereo sensors, stemming from incorrect depth priors due to the triangulation of flying pixels, which the MM-2FSK method cannot correct.
Figure 8. Top-down views of the RF ToF reconstructions obtained with the MM-2FSK method, fused with the ground-truth MVS point cloud. From left to right, we vary the supporting sensor providing the depth prior. From top to bottom, we display the Flowerpot (Transparent), V2 Metal Plate, and Bunny objects.
Depth Prior $\text{C}_{\text{g}}$ $\text{C}_{\text{s}}$ P $\text{P}_{\text{e}}$
MVS 0.51 0.18 0.19 0.17
NIR ToF 1.19 1.67 1.58 1.48
Active Stereo 0.82 1.74 1.36 1.36
Passive Stereo 0.92 1.91 1.29 1.41
Table 5. Ablation study of the MM-2FSK method with different depth priors, each from another optical depth imager. The mean depth deviation from the ground truth, given in centimeters, is averaged over all objects at 30 cm distance. The best and second best results per metric are highlighted.
Moreover, Table 5 lists the mean depth deviation of all objects in relation to the varying depth priors. By assessing the depth deviation always relative to the MVS setup, we expect that its respective depth prior yields the best performance.
Consistent with earlier evaluations (cf. Section 6.2.1), no single sensor outperforms others across all objects. For 3D errors ($\text{C}_{\text{g}}$and $\text{C}_{\text{s}}$), the active stereo and NIR sensors yield the best performance, while for projective errors (Pand $\text{P}_{\text{e}}$), both passive stereo and active stereo sensors provide the most accurate results.
9. Limitations
During sensor characterization, we observed that, even though the depth deviation of the MIMO imaging radar is on par with that of optical sensors, the reconstructions exhibit considerably more holes, where no valid depth is estimated. While it is intuitive to assume that reconstruction quality is influenced by object material (limiting the returned signal amount — akin to optically transmissive materials), we observe that in our experiments the object geometry is the primary factor of influence and has a greater impact than in optical sensors; however, disentangling the effects of geometry and material remains challenging, as precise impacts on fine-grained surface details cannot be easily assessed. Without a direct mapping between point targets and RX antennas, depth evaluation concerning these surface-level details is infeasible without backprojection, or any other depth processing algorithm. On the other hand, the reconstructed outcomes after signal processing may not align with reality, as these methods typically incorporate a systematic bias by relying on the Born approximation (Ahmed, 2014). To overcome these limitations, we highlighted one particular work of Hofmann et al. (2025), taking the first step towards automatic material characterization in the radio-frequency domain using inverse rendering. We anticipate that further analysis of these material parameters, combined with improvements of radio-frequency simulation frameworks (Schüßler et al., 2021), will considerably aid in disentangling the potential error sources behind missing reconstructions and enhancing current signal processing methods.
The proposed evaluation framework for sensor characterization is tailored to point cloud comparisons, and is, therefore, independent of the RF signal processing algorithm; however, it requires the spatial co-localization of sensors. To achieve the latter, it needs to be verified, whether the respective spatial calibration method may be applicable to other high-resolution radar systems; alternatively, it can be substituted with any other calibration method tailored to the radar system of interest.
Furthermore, the object reconstructions were evaluated solely for valid locations in the ground-truth data, excluding artifacts like ghost targets or other forms of noise that may arise from violations of the Born approximation. Lastly, we did not capture different orientations of flat objects, which would be an interesting future direction to investigate object orientation in isolation from geometry complexity.
10. Conclusion
We presented a novel multimodal dataset, MAROON, that allows us to characterize, for the first time, near-field MIMO imaging radars in direct relation with traditional depth imagers from the optical frequency domain for close-range applications. The dataset comprises depth images of a variety of objects, synchronously captured by four mutually calibrated depth imagers and a ground-truth multi-view stereo system. We subsequently analyzed the data within a comprehensive evaluation framework, offering quantitative and qualitative perspectives on each sensor’s depth deviation across multiple metric types, objects, and object-to-sensor distances. The findings presented are based on aggregate trends and individual object analyses that contribute to the understanding of the addressed sensor characteristics; however, we believe that our dataset still invites further analysis, exploiting the high diversity of the 45 objects that could not be fully addressed in the scope of this paper.
Moreover, we presented two exemplary applications, utilizing the collected data. First, we built upon previous work to characterize the materials of our captured objects, which is an interesting future direction to disentangle material-specific effects from geometric influences. Second, we conducted extended experiments on a recently proposed multimodal depth estimation approach (Wirth et al., 2025), using our dataset as a baseline to evaluate its performance. In connection with this work, we examined the impact of different optical sensor modalities to identify suitable depth priors for radar signal processing.
We hope that by highlighting these promising research directions, along with the release of our MAROON dataset, our work will give rise to further study of multimodal sensor systems in a joint reference frame.